OpenStudy (anonymous):

Which of the following is equivalent to log 6 of 21 ? log 21 log 1/6 log 21/log 6 log 6/ log 21

OpenStudy (anonymous):

a

OpenStudy (anonymous):

actually it is C as always

OpenStudy (anonymous):

by the change of base formula \[\log_6(21)=\frac{\log(21)}{\log(6)}\]

OpenStudy (anonymous):

i disagree